My son is doing Algebra foundations. This is the problem that he needs to solve and I think I am solving it correctly however, I need to make sure. The speed (s) that Brandon can ride his bike if he rides 3/5 of an hour and travels 4 miles is given by the equation 4 = 3/5 (s) What is Brandon's speed?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

The speed (s) that Brandon can ride his bike if he rides 3/5 of an hour and travels 4 miles is given by the equation 4 = 3/5 (s) What is Brandon's speed?
----
rate = distance/time = 4/(3/5) = 20/3 = 6 2/3 mph
=====================
Cheers,
Stan H.

Answer by solver91311(24713)



Distance equals rate times time, so yes, if the distance is 4 miles, the time is , and the rate is , then is correct. Now all you need to do is solve the equation for -- multiply both sides by .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism